In a computer lab, there are 3 computers for every 6 students. How many computers will be needed for 24 students?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that,
Number of computer required for 6 students = 3
So, number of computer required for 1 student = (3/6)
=1⁄2
So, number of computer required for 24 students = 24 × 1⁄2
= 24/2
= 12
∴Number of computer required for 24 students is 12 computers.

Thanks